# ScribblePrompt: Fast and Flexible Interactive Segmentation for Any Medical Image

### v1

@@ Appendix C User Study
Study Design. The goal of the user study was to compare ScribblePrompt to the best click-focused baseline method in terms of accuracy (Dice of the final segmentations), efficiency (time to achieve the desired segmentations) and user experience (perceived effort). Participants were given time to familiarize themselves with both models on a fixed set of practice images. Afterwards they used each model to segment a series of nine test images from nine tasks (Figure 15).
The order in which participants used the models, and which image the users were assigned to segment with each model for each task was randomized. We randomly selected one training image per task to include in the set of practice images. We randomly selected two test images per task from the evaluation sets, for a total of 18 images used in the study. The models were also annonymized (i.e., “Model A” and “Model B”). We informed participants that one model was designed to be used with clicks and bounding boxes, while the other was designed for use with clicks, bounding boxes, and scribbles.
@@
We had a total of 29 participants with 16 participants completing all of the segmentations and the exit survey. We observed a higher attrition rate among participants who were assigned to use SAM first, even after being able to freely try out both models during the “practice” phase. Among the 13 participants assigned to use SAM first, 62% did not finish all of their segmentations, compared with 31% among the 16 assigned to use ScribblePrompt first. We report results on the 16 participants who completed both the segmentations and the exit survey.
Implementation. Each participant used a web-based interface powered by a Nvidia Quatro RTX8000 GPU with 4 CPUs. The interface was developed in Python using the Gradio library [2]. The interface had a “practice” mode in which users could freely switch between the two models and images from the set of practice images. After experimenting with both models, users clicked a button to begin “recorded activity” mode in which users were led through performing specific segmentation tasks with specific models. Users provided positive/negative scribble inputs, positive/negative click inputs and/or bounding box inputs, and then clicked a button to receive a prediction from the model.
@@
Visualizations. We visualize some of the interactions used by study participants and the resulting predictions in Figure 15. Study participants used denser clicks when prompting SAM compared to when prompting ScribblePrompt-Unet.
Appendix D Experiments
D.1 Implementation Details
Baselines. For baseline methods using the SAM architecture, we evaluate the models in both “single mask” and “multi-mask” mode. For each baseline method and interaction procedure, we selected the best performing mode based on the average Dice across the validation data and report results on test data using that mode. In the results with simulated clicks and scribbles by dataset (Figures 19, 20, 21, 22, 23, and 24), we show results using both modes.
For baseline methods using the SAM architecture, which was not designed to take scribble masks as input, we convert the scribble masks to sets of positive and negative clicks for every non-zero pixel in the scribble masks.
D.2 Simulated Bounding Boxes
We evaluate models with simulated bounding box prompts.
Setup. We evaluate segmentation accuracy using Dice score after a single bounding box prompt. We simulate bounding boxes using the same procedure as used during training ScribblePrompt: we find the minimum enclosing bounding box for the ground truth label and then enlarge each dimension by $r\sim U[0,20]$ pixels to account for human error.
In addition to SAM [61] and SAM-Med2d [25], we also include MedSAM [82] as a baseline. MedSAM is a SAM model with ViT-b backbone that was fine-tuned on a collection of over 1M medical images [^2] MedSAM used two of our evaluation datasets (TotalSegmentator and BUID) as training datasets. with one bounding box input per example. We exclude MIDeepSeg [80] from this evaluation because it is not designed to make predictions from a single bounding box input.
For methods using the SAM architecture, we apply the pixel normalization scheme in [61] to images before inference. Upon further investigation, MedSAM performed better with images rescaled to $[0,1]$; we report results for MedSAM with both normalization schemes.
Results. Figure 16 shows mean Dice after one bounding box prompt. Figure 17 shows results by dataset. ScribblePrompt-SAM has the highest Dice on average after one bounding box prompt.
Figure 16: Results with simulated bounding boxes. Mean Dice on test data from 12 datasets with one simulated bounding box prompt, weighting each dataset equally. $*{}^{*}$ indicates MedSAM with input images re-scaled to $[0,1]$ instead of the pixel normalization from [61]. Errorbars show 95% CI.
Figure 17: Results with simulated bounding boxes by dataset. Mean Dice after one simulated bounding box prompt. Among the evaluation datasets, bounding box prompts are the most effective for BUID, a breast ultrasound dataset. $*{}^{*}$ indicates MedSAM with input images re-scaled to $[0,1]$ instead of the pixel normalization from [61]. Errorbars show 95% CI.
D.3 Simulated Scribbles and Clicks
We provide more detailed results for experiments with simulated scribbles and clicks presented in the main paper.
Setup. See Section 5.2 in the main paper.
Results. Figure 18 shows Dice vs. steps of interaction for three simulated click-focused procedures and three simulated scribble-focused procedures. On average, ScribblePrompt-Unet and ScribblePrompt-SAM have the highest Dice at all steps for all of the simulated interaction procedures.
Figures 19, 20, and 21 show results by dataset for scribble-focused interaction procedures. Figures 22, 23, and 24 show results by dataset for the click-focused interaction procedures.
Figure 18: Results with simulated clicks and scribbles. We evaluate methods using three simulated scribble procedures and three simulated click procedures. We measure Dice averaged across twelve evaluation sets (the test splits of the nine validation and three test datasets), weighting each dataset equally. Shaded regions show 95% CI from bootstrapping.
Appendix E Inference Runtime
@@
Setup. We measure inference time for a random input with a scribble covering 128 pixels. We report mean and standard deviation of inference time across 1,000 runs.
@@
Results. We show performance results on a Nvidia RTX8000 GPU in Table 4. ScribblePrompt-Unet is faster than the baseline methods on a GPU. Unlike the transformer-based models, the inference time for each prediction is constant with the number of input interactions for ScribblePrompt-Unet.
@@
Table 4: GPU Performance Summary. We report the mean and standard deviation for inference time on a Nvidia RTX8000 GPU for a prediction with a scribble input covering 128 pixels. ScribblePrompt SAM has the same runtime and number of parameters as SAM ViT-b.
Model Param. Runtime (ms)
SAM (ViT-h) 641M $504.36\pm 57.72$
SAM (ViT-b) 94M $133.85\pm 24.26$
MIDeepSeg 3M $65.75\pm 21.87$
SAM-Med2d w/ adapter 271M $35.06\pm 12.88$
SAM-Med2d w/o adapter 91M $24.86\pm 9.56$
ScribblePrompt-Unet 4M $1.96\pm 0.20$
Appendix F Data
We build on MegaMedical [20] to include a total of 77 datasets. We added datasets with a particular focus on Microscopy, X-Ray, and Ultrasound modalities, which were not well represented in the original MegaMedical. The full list of datasets is provided in Tables 6 and 7 and 8.
We define a 2D segmentation task as a combination of (sub)dataset, axis (for 3D modalities), and label. For datasets with multiple segmentation labels, we consider each label separately as a binary segmentation task and for 3D modalities we use the slice with maximum label area and the middle slice for each subject. For multi-annotator datasets, we treated each annotator as a separate label. For instance segmentation datasets, we sampled one instance at a time during training. All images were resized to $128\times 128$ and re-scaled to $[0,1]$. For methods using the SAM image encoder, we convert images to RGB and apply the pixel normalization scheme in [61].
The division of data for training, validation and evaluation is summarized in 5. All 77 datasets were split into 60% train, 20% validation, and 20% test, although some splits were not used. The 77 datasets were divided into 65 training datasets (Tables 6 and 7), 9 validation datasets (ACDC, BTCV Cervix, BUID, DRIVE, HipXRay, PanDental, SCD, SpineWeb, WBC) and 3 test datasets (COBRE, SCR, TotalSegmentator). We used the “train” splits from the 65 training datasets to train the models. We use the “validation” splits from the 9 validation datasets during model development for model selection.
We report final evaluation results across 12 evaluations sets consisting of the “test” splits of the 9 validation datasets and “test” splits of the 3 test datasets to maximize the diversity of tasks and modalities in our evaluation set. No data from the 9 validation datasets or 3 test datasets were seen by the models during training.
Development Phase
Datasets Training Validation Evaluation
Training (n=65)
Validation (n=9)
Test (n=3)
Table 5: Dataset split overview. Datasets were split into 60% train, 20% validation and 20% test. Data from the “train” splits of the 65 training datasets were used to train the models. The models did not see any data from the validation or test datasets during training. Data from the “validation” split of the 9 validation datasets was used for model selection. We report final results on 12 “evaluation sets”: data from the “test” splits of the 9 validation datasets and the “test” splits of the 3 test datasets.
Dataset Name Description # of Scans Image Modalities
AbdominalUS [127]
Abdominal organ segmentation
1,543
Ultrasound
AMOS [54]
Abdominal organ segmentation
240
CT, MRI
BBBC003 [77]
Mouse embryos
15
Microscopy
BBBC038 [21]
Nuclei instance segmentation
670
Microscopy
BrainDevelopment [39, 38, 64, 115]
Adult and neonatal brain atlases
53
Multimodal MRI
BrainMetShare[41]
Brain tumors
420
Multimodal MRI
BRATS [10, 11, 93]
Brain tumors
6,096
Multimodal MRI
BTCV Abdominal [66]
13 abdominal organs
30
CT
BUSIS [146]
Breast tumors
163
Ultrasound
CAMUS [67]
Four-chamber and Apical two-chamber heart
500
Ultrasound
CDemris [55]
Human left atrial wall
60
CMR
CHAOS [57, 56]
Abdominal organs (liver, kidneys, spleen)
40
CT, T2-weighted MRI
CheXplanation [112]
Chest X-Ray observations
170
X-Ray
CT2US [121]
Liver segmentation in synthetic ultrasound
4,586
Ultrasound
CT-ORG[107]
Abdominal organ segmentation (overlap with LiTS)
140
CT
DDTI [99]
Thyroid segmentation
472
Ultrasound
EOphtha [30]
Eye microaneurysms and diabetic retinopathy
102
Optical camera
FeTA [98]
Fetal brain structures
80
Fetal MRI
FetoPlac [12]
Placenta vessel
6
Fetoscopic optical camera
FLARE [81]
Abdominal organs (liver, kidney, spleen, pancreas)
361
CT
HaN-Seg [101]
Head and neck organs at risk
84
CT, T1-weighted MRI
HMC-QU [31, 60]
4-chamber (A4C) and apical 2-chamber (A2C) left wall
292
Ultrasound
I2CVB [68]
Prostate (peripheral zone, central gland)
19
T2-weighted MRI
IDRID [102]
Diabetic retinopathy
54
Optical camera
ISBI-EM [22]
Neuronal structures in electron microscopy
30
Microscopy
ISIC [26]
Demoscopic lesions
2,000
Dermatology
ISLES [46]
Ischemic stroke lesion
180
Multimodal MRI
KiTS [45]
Kidney and kidney tumor
210
CT
LGGFlair [19, 89]
TCIA lower-grade glioma brain tumor
110
MRI
LiTS [15]
Liver tumor
131
CT
LUNA [116]
Lungs
888
CT
MCIC [37]
Multi-site brain regions of schizophrenic patients
390
T1-weighted MRI
MMOTU [147]
Ovarian tumors
1,140
Ultrasound
MSD [118]
Large-scale collection of 10 medical segmentation datasets
3,225
CT, Multimodal MRI
MuscleUS [88]
Muscle segmentation (biceps and lower leg)
8,169
Ultrasound
NCI-ISBI [16]
Prostate
30
T2-weighted MRI
NerveUS [95]
Nerve segmentation
5,635
Ultrasound
OASIS [47, 86]
Brain anatomy
414
T1-weighted MRI
OCTA500 [69]
Retinal vascular
500
OCT/OCTA
PanNuke [35]
Nuclei instance segmentation
7,901
Microscopy
PAXRay [114]
92 labels covering lungs, mediastinum, bones, and sub-diaphram in Chest X-Ray
852
X-Ray
Dataset Name Description # of Scans Image Modalities
PROMISE12 [74]
Prostate
37
T2-weighted MRI
PPMI [87, 29]
Brain regions of Parkinson patients
1,130
T1-weighted MRI
QUBIQ [92]
Collection of 4 multi-annotator datasets (brain, kidney, pancreas and prostate)
209
T1-weighted MRI, Multimodal MRI, CT
ROSE [83]
Retinal vessel
117
OCT/OCTA
SegTHOR [65]
Thoracic organs (heart, trachea, esophagus)
40
CT
SegThy [62]
Thyroid and neck segmentation
532
MRI, Ultrasound
ssTEM [36]
Neuron membranes, mitochondria, synapses and extracellular space
20
Microscopy
STARE [48]
Blood vessels in retinal images (multi-annotator)
20
OCT
ToothSeg [51]
Individual teeth
598
X-Ray
VerSe [78]
Individual vertebrae
55
CT
WMH [63]
White matter hyper-intensities
60
Multimodal MRI
WORD [79]
Abdominal organ segmentation
120
CT
Table 8: Validation and test datasets. We assembled the following set of datasets to evaluate ScribblePrompt and baseline methods. For the relative size of datasets, we include the number of unique scans (subject and modality pairs) that each dataset has.
Dataset Name Description # of Scans Image Modalities
ACDC [14]
Left and right ventricular endocardium
99
cine-MRI
BTCV Cervix [66]
Bladder, uterus, rectum, small bowel
30
CT
BUID [5]
Breast tumors
647
Ultrasound
COBRE [4]
Brain anatomy
258
T1-weighted MRI
DRIVE [122]
Blood vessels in retinal images
20
OCT
HipXRay [42]
Ilium and femur
140
X-Ray
PanDental [1]
Mandible and teeth
215
X-Ray
SCD [103]
Sunnybrook Cardiac Multi-Dataset Collection
100
cine-MRI
SCR [125]
Lungs, heart, and clavicles
247
X-Ray
SpineWeb [148]
Vertebrae
15
T2-weighted MRI
TotalSegmentator [134]
104 anatomic structures (27 organs, 59 bones, 10 muscles, and 8 vessels)
1,204
CT
WBC [149]
White blood cell cytoplasm and nucleus
400
Microscopy
Figure 19: Results per dataset with line scribbles. During the first step we simulate three positive line scribbles and three negative line scribbles. In subsequent steps, we simulate one (positive or negative) correction line scribble based on the error region between the previous prediction and ground truth segmentation. Each line scribble covers a maximum of 128 pixels.
Figure 20: Results per dataset with centerline scribbles. During the first step, we simulate one positive and one negative centerline scribble. In subsequent steps, we simulate one (positive or negative) correction centerline scribbles based on the error region region between the previous prediction and ground truth segmentation. Each centerline scribble covers a maximum of 128 pixels.
Figure 21: Results per dataset with contour scribbles. During the first step, we simulate one positive and one negative contour scribble based on the ground truth label. In subsequent steps, we simulate one (positive or negative) correction contour scribble based on the error region region between the previous prediction and ground truth segmentation.Each contour scribble covers a maximum of 128 pixels.
Figure 22: Results per dataset with center clicks. During the first step, one positive click is placed at the center of the largest component of the ground truth segmentation. In subsequent iterations, one (positive or negative) correction click is placed at the center of the largest component of the error region between the previous prediction and ground truth segmentation.
Figure 23: Results per dataset with random clicks. During the first step, one positive click is randomly sampled from the ground truth segmentation. In subsequent steps, one (positive or negative) correction click is randomly sampled from the error region between the previous prediction and ground truth segmentation.
Figure 24: Results with random warm start click procedure. During the first step, three positive clicks are randomly sampled from the ground truth segmentation and three negative clicks are randomly sampled from the background. In subsequent steps, one (positive or negative) correction click is placed at the center of the largest component of the error region between the previous prediction and ground truth segmentation.

### v2

@@ Appendix G User Study
Study Design. The goal of the user study was to compare ScribblePrompt to the best click-focused baseline method in terms of accuracy (Dice of the final segmentations), efficiency (time to achieve the desired segmentations) and user experience (perceived effort). Participants were given time to familiarize themselves with both models on a fixed set of practice images. Afterwards they used each model to segment a series of nine new test images from nine tasks that were not seen by the model during training (Fig. 35).
The order in which participants used the models, and which image the users were assigned to segment with each model for each task was randomized. We randomly selected one training image per task to include in the set of practice images. We randomly selected two test images per task and randomized the assignment of each image to each model for each participants. Each participant segmented a total of 18 images during the study. The models were also annonymized (i.e., “Model A” and “Model B”). We informed participants that one model was designed to be used with clicks and bounding boxes, while the other was designed for use with clicks, bounding boxes, and scribbles.
@@
We had a total of 29 participants with 16 participants completing all of the segmentations and the exit survey. We observed a higher attrition rate among participants who were assigned to use SAM first, even after being able to freely try out both models during the “practice” phase. Among the 13 participants assigned to use SAM first, 62% did not finish all of their segmentations, compared with 31% among the 16 assigned to use ScribblePrompt first. We report results on the 16 participants who completed all the segmentations and the exit survey.
Implementation. Each participant used a web-based interface powered by a Nvidia Quatro RTX8000 GPU with 4 CPUs. Participants segmented the images at $256\times 256$ resolution. The interface was developed in Python using the Gradio library [2]. The interface had a “practice” mode in which users could freely switch between the two models and images from the set of practice images. After experimenting with both models, users clicked a button to begin “recorded activity” mode in which users were led through performing specific segmentation tasks with specific models. Users provided positive/negative scribble inputs, positive/negative click inputs and/or bounding box inputs, and then clicked a button to receive a prediction from the model.
@@
Visualizations. We visualize some of the interactions used by study participants and the resulting predictions in Fig. 35. Study participants used denser clicks when prompting SAM compared to when prompting ScribblePrompt-UNet.
Appendix H Inference Runtime
@@
Setup. We measure inference time for a random input with a scribble covering 128 pixels. We report mean and standard deviation of inference time across 1,000 runs on a single CPU and on a Nvidia Quatro RTX8000 GPU.
@@
Results. We show performance results in Tab. 8. On a single CPU, ScribblePrompt-UNet requires $0.27\pm 0.04$ sec per prediction, enabling the model to be used in real-time even in low-resource environments. ScribblePrompt-UNet is faster than the baseline methods on a GPU. Unlike the transformer-based models, the inference time for each prediction is does not depend on the number of of input interactions for ScribblePrompt-UNet.
@@
Table 8: Performance Summary. We measure inference time separately on a single CPU and on an Nvidia Quatro RTX8000 GPU for a prediction with a random scribble input covering 128 pixels. We report mean and standard deviation across 1,000 runs. ScribblePrompt-SAM and MedSAM use the same architecture as SAM ViT-b. Best and second best are highlighted.
Architecture   Param.   CPU Runtime (sec)   GPU Runtime (ms)
SAM (ViT-h) 641M $130.79\pm 7.96$ $504.36\pm 57.72$
SAM (ViT-b) 94M $13.59\pm 0.77$ $133.85\pm 24.26$
SAM-Med2D w/ adapter 271M $1.23\pm 0.07$ $35.06\pm 12.88$
SAM-Med2D w.o. adapter 91M $0.63\pm 0.02$ $24.86\pm 9.56$
MIDeepSeg 3M $0.08\pm 0.02$ $65.75\pm 21.87$
ScribblePrompt-UNet 4M $0.27\pm 0.04$ $1.96\pm 0.20$
Appendix I Ablations
We conduct two ablations of important ScribblePrompt design decisions: (1) synthetic label inputs used during training, and (2) types of prompts simulated during training. We report results on the validation splits of nine validation datasets that were unseen during training.
I.1 Synthetic Labels
Setup. We trained ScribblePrompt-UNet and ScribblePrompt-SAM with different values of $p_{synth}$, the probability of sampling a synthetic label.
Results. Training with some synthetic labels improves both ScribblePrompt-UNet and ScribblePrompt-SAM’s performance on validation data from nine (validation) datasets not seen during training (Fig. 36, 37). For both ScribblePrompt-UNet and ScribblePrompt-SAM, training with 50% synthetic labels leads to the highest Dice on unseen datasets at inference time.
Figure 36: Probability of synthetic labels during training for ScribblePrompt-UNet. We report change in Dice relative to ScribblePrompt-UNet trained without any synthetic labels ($p_{synth}=0$). We show Dice after five steps of simulated interactions following six different (inference-time) interaction procedures. Errorbars show 95% CI.
Figure 37: Probability of synthetic labels during training for ScribblePrompt-SAM. We report change in Dice relative to ScribblePrompt-SAM trained without any synthetic labels ($p_{synth}=0$). We show Dice after five steps of simulated interactions following six different (inference-time) interaction procedures. Errorbars show 95% CI.
I.2 Prompt Types
Setup. We evaluate ScribblePrompt-UNet models trained with different combinations of prompts, compared to the complete ScribblePrompt-UNet:
  • ScribblePrompt-UNet (scribbles) trained on boxes and scribbles.
  • ScribblePrompt-UNet (clicks) trained on boxes and clicks.
  • ScribblePrompt-UNet (random clicks) trained on boxes and random clicks.
Results. Fig. 38 shows results for six different inference-time interaction procedures. ScribblePrompt-UNet trained with scribbles, clicks, and bounding boxes predicts segmentations more accurately than do ablated versions of ScribblePrompt-UNet.
Figure 38: Ablation of interactions during training. We report Dice after five steps of simulated interactions following six inference-time interaction procedures. Error bars show 95% CI from bootstrapping.